OpenStudy (anonymous):

the expressiion cos^2A-cos2A is equivalent to?

Directrix (directrix):

sin^2 A or more clearly written as (sin A ) ^2

OpenStudy (anonymous):

cos squared a

Directrix (directrix):

Is that the answer you computed? I need to check my work.

OpenStudy (anonymous):

i have no answer. i have a question cos^2 A - Cos2A

Directrix (directrix):

I didn't follow when you wrote "cos squared a" just now. I thought you were saying that is the answer.

OpenStudy (anonymous):

oh no. sorry

Directrix (directrix):

Is cos squared a another problem to work or simplify?

OpenStudy (anonymous):

no i just have one question

OpenStudy (anonymous):

the expression cos squared a - cos 2 a is equivalent to what

OpenStudy (anonymous):

i have 4 answer choices

OpenStudy (anonymous):

sin squared a -sin squared a cos squared a + 1 -cos squared a - 1

Directrix (directrix):

sin squared a is the answer.

OpenStudy (anonymous):

can you explain please?

OpenStudy (lgbasallote):

cos2A = cos^2A - sin^2A cos^2A -(cos^2A - sin^A) the answer is sin^2A

Directrix (directrix):

cos^2A-cos2A = (1 - sin² A) - [ 1 - 2 sin^2(x) ] = 1 - sin² A - 1 + 2 sin² (x) = sin² (x) ==> Answer Notes: sin^2(x) + cos^2(x) = 1 --> Identities to memorize cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = 1 - 2 sin^2(x) http://www.math.com/tables/trig/identities.htm